A 90.00 g sample of Na₂CO₃⋅xH2O is heated until all of the water is driven off, forming 0.314 mol of the anhydrous salt, Na₂CO₃. The molar mass of Na₂CO₃ is 105.99 g/mol. What is the percent by mass of H2O in the hydrated compound, Na₂CO₃⋅xH2O?


1
Expert's answer
2020-11-09T14:10:35-0500

Solution:

The chemical reaction equation:

Na2CO3⋅xH2O --> Na2CO3 + xH2O


According to the law of conservation of mass, the mass of the products in a chemical reaction must equal the mass of the reactants.

or

m(Na2CO3⋅xH2O) = m(Na2CO3) + m(xH2O)


m(Na2CO3) = Moles of Na2CO3 × Molar mass of Na2CO3

m(Na2CO3) = 0.314 mol × 105.99 g/mol = 33.28 g


m(xH2O) = m(Na2CO3⋅xH2O) - m(Na2CO3) = 90.00 g - 33.28 g = 56.72 g


Mass percent of a component = (mass of component / total mass of sample) × 100%

The percent by mass of H2O in the hydrated compound (Na2CO3⋅xH2O) is:

w(H2O) = (m(xH2O) / m(Na2CO3⋅xH2O)) × 100%

w(H2O) = (56.72 g / 90.00 g) × 100% = 63.02%

w(H2O) = 63.02%


Answer: The percent by mass of H2O in the hydrated compound is 63.02%.